Bob invests $2000 at 6% interest for 5 years. How much will he earn in 5 years? How much will he have after 5 years?
Answer by fcabanski(1391)

If the problem doesn't specify, then it's simple interest.


The formula for simple interest


P = Principal = 2000

r = interest rate =6% = 6/100 = .06

t = time (in years) = 5

I = interest earned = solution

I=P*r*t


I =2000*.06*5= $600


He will earn $600 in 5 years. He will have the original principal plus the interest after 5 years.


$2000+$600=$2600.

I provide online tutoring ($30/hr) and personal problem solving ($3.50-$5.50 per problem. Contact me (check the profile "website" which is my email address.